Assume that x and y are differentiable functions of t and
y^2 = 2xy + 24. Find dx/dt given that x = 1, y = 6, and dy/dt = 2

Answer by Alan3354(69443)

Assume that x and y are differentiable functions of t and
y^2 = 2xy + 24. Find dx/dt given that x = 1, y = 6, and dy/dt = 2
-------
y^2 = 2xy + 24
y^2/2 = xy + 12
y(dy/dt) = x*dy/dt + y*dx/dt
6*2 = 1*2 + 6*dx/dt
dx/dt = 5/3